The Krewe of Orpheus maintains a supply of swizzle sticks for events throughout the year. Demand for swizzle sticks is shockingly low, a quick check of krewe records from last year reveals that they used only 585,000, but the krewe president believes that they should be good stewards of what they have, so they seek to manage this inventory using the EOQ policy, although they prefer to refer to it as an EOKrewe policy for obvious reasons. Swizzle sticks are not expensive items, they cost a nickel apiece largely due to the club logo printed on each one. This also serves to increase the lead time as they can't be obtained from a standard restaurant supply house. Instead, they must be ordered with an eye towards the six-day lead time. It costs $15 to place an order, most of this cost is a result of explaining the meaning of "Laissez les bons temps rouler" and why it should be printed on the edge of each swizzle stick. Holding cost is 20% of purchase price.

A) What would the order quantity for swizzle sticks be?

B) What is the average inventory if they order at their EOKrewe quantity?

C) What is their total inventory policy cost (excluding cost of goods) if they order at their EOKrewe amount?
Answer #1

Given in the case

The Annual demand (D) for sticks = 585,000 (Given)

Ordering cost per order (Co ) = $15 (Given)

Holding Cost = 20% of purchase price

The Annual holding cost (H) for sticks = 20% * cost per piece

= 20% * 0.05 ( a nickel = 5 cents) = 0.01 per piece

Annual holding cost (H) = 0.01 per piece

Ordering Cost (S) = 15$

\small EOQ =\sqrt{}(2DS)/H

\small = \sqrt{}(2\times 585,000\times 15)/0.01

= 41,892.72 pieces

EOQ = 41,893 pieces (Approximate)

The Average inventory level = (0 + Q)/2 = Q/2

= ( 0 + 41,893)/2 = 20946.5

The Average inventory level = 20947 Pieces ( Approximate)

Total ordering cost = (D/Q) × Co

= (585,000 / 41,893) × 15

Total ordering cost = $209.76

The Total no. of production runs that required per year = Total demand / EOQ

= 585,000 / 41,893

Total no. of production runs that required per year = 14